hi, my name is sherry....please help me. I really hate word problems.
A bank invested $480,000 last year in mutual funds. th consevativefund earned 8% simpl interest. the growth fund earned 12% simple interest.At he end of theyear , the bank had earned $49,000 from these mutual funds. how much did it invest in each fund? here is what i got for the first part
what is the amount invested at 12% in growth fund?answer; 5880
what is the amount invested at 8% in the conservative fund? answer:3920
i divided each question by 49,ooo and use the percent in te question. is this right? thank you so much Answer by rfer(16322) 
 Ok Sherry here we go.
It is necessary to use the invested amount in the calculation, not the interest earned so we get this relationship between the two funds. The relationship equals the interest earned.
.08x+.12(480000-x)=49000
.08x+57600-.12x=49000
-.04x=-8600
x=$215,000
480000-x=$265,000